the probability that a network will be shut down on any given day is .0003. What is the probability that the network will be shut down in 5000 days.
Answer by stanbon(75887)

the probability that a network will be shut down on any given day is .0003. What is the probability that the network will be shut down in 5000 days.
-----
P(down >= 1 in 5000 days) = 1 - P(up for 5000 days)
---
= 1 - 0.9997^5000
---
= 1 - 0.2231
---
= 0.7769
==================
Cheers,
Stan H.